OpenStudy (anonymous):

An olympic size rectangular swimming pool has dimentions 100m by 50m. calculate the length of the diagonal of the pool as well as the angle the diagonal makes with the length

OpenStudy (anonymous):

a^2 + b^2 = c^2 where a=100, b=50, and c is the diagonal that you are looking for. Solve for c.

OpenStudy (anonymous):

what does the 2 stand for

OpenStudy (anonymous):

Also, sin x = 50/100, where x is the angle you are looking for. The 2's are exponents. Rewriting:\[a ^{2} + b ^{2} = c ^{2}\]

OpenStudy (anonymous):

so would it be right if i said \[\sin \theta =\frac{ 50 }{ 100 }\]

OpenStudy (anonymous):

Solving for c: \[c = \sqrt{a ^{2} + b ^{2}}\]. Yes, you could use theta instead of x for the angle.

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

You're welcome! Good luck in all your studies!